# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v6

@@ 3.1 DRAM Refresh
DDR5 introduces Same Bank Refresh (SBR), which refreshes one bank in each bank group at a time instead of refreshing all banks as in DDR4 [78, 83]. SBR improves bank availability as the MC can access the non-refreshing banks while certain banks are being refreshed. DDR5 implements SBR with a new REFsb command [72]. Implementing REFsb necessitates changes in the DRAM interface and MC.
@@ 3.2 RowHammer Protection
In DDR4, DRAM vendors implement in-DRAM RowHammer protection mechanisms by performing Targeted Row Refresh (TRR) within the slack time available when performing regular refresh. However, prior works have shown that in-DRAM TRR is vulnerable to certain memory access patterns [74, 75, 76]. DDR5 specifies the Refresh Management (RFM) that an MC implements to aid in-DRAM RowHammer protection. As part of RFM, the MC uses counters to keep track of row activations to each DRAM bank. When a counter reaches a specified threshold value, the MC issues the new RFM command to DRAM chips. A DRAM chip then performs an undocumented operation with the time allowed for RFM to mitigate RowHammer. Since the first RowHammer work [10] in 2014, it took about 7 years to introduce DDR5 RFM, which requires significant changes in the DRAM interface and the MC design. Still, RFM is likely not a definitive solution for RowHammer as it does not outline a RowHammer defense with security proof. Rather, RFM provides additional time to a DRAM chip for internally mitigating the RowHammer.
@@ 3.3 Memory Scrubbing
DDR5 adds support for on-die ECC and in-DRAM scrubbing. A DDR5 chip internally performs ECC encoding and decoding when the chip is accessed. To perform DRAM scrubbing, the MC must periodically issue the new scrub command (for manual scrubbing) or all-bank refresh command (for automatic scrubbing). Similar to Same Bank Refresh and RFM, enabling in-DRAM scrubbing necessitated changes in the DRAM interface and in the MC design.
4 Self-Managing DRAM
To enable autonomously and efficiently performing maintenance operations in a DRAM chip, we introduce the Self-Managing DRAM (SMD) architecture. SMD is part of a continuum between fully master-slave and fully request-reply interfaces. It tries to achieve a balance that enables faster innovation and easier adoption of new ideas.
@@ 4.1 Overview of SMD
Organization. SMD preserves the general DRAM interface and uses a single uni-directional pin (which already exists in DDR4/5, see section 7.4) in the physical interface of a DRAM chip. This pin is used for transmitting the ACT_NACK signal from the DRAM chip to the MC.
Fig. 2 shows the organization of a bank in an SMD chip. SMD divides a DRAM bank into multiple lock regions of equal size. SMD stores an entry in a per-bank Lock Region Table (LRT) to indicate whether or not a lock region is reserved for performing a maintenance operation.
Operation. A maintenance operation takes place as the Lock Controller sets the lock bit in the LRT entry that corresponds to the lock region on which the maintenance operation is to be performed. When the MC attempts to open a row in a locked region, the Lock Controller generates an ACT_NACK.
A maintenance operation and an ACT can be performed concurrently in the same bank on different lock regions.333An ECC scrubbing operation (section 5.3) uses the data bus, which is a shared resource across all regions. Therefore, a scrubbing operation reservers all lock regions inside a bank at once to prevent any other maintenance operation or memory access from happening at the same time. To enable this, SMD implements a row address latch in order to drive two local row decoders with two independent row addresses.444One row address latch per lock region enables two or more maintenance operations to concurrently happen on different lock regions within a bank. To keep our design simple, we restrict a maintenance operation to happen in one lock region while the MC accesses another lock region. Our design can be easily extended to increase concurrency of maintenance operations across multiple lock regions. When a maintenance operation and an ACT arrive at the same lock region at the same time, the MUX prioritizes the maintenance operation and the SMD chip rejects the ACT.
@@ 4.2 Region Locking Mechanism
A maintenance operation can be performed on a lock region only after locking it. Therefore, a maintenance mechanism must lock the region that includes the rows that should undergo a maintenance. A maintenance mechanism can only lock a region that is not already locked, which prevents different maintenance mechanisms from interfering with each other.
Lock Region Size. A lock region consists of a fixed number of consecutively-addressed DRAM rows. To simplify the design, we set the lock region size to one or multiple subarrays. This is because a maintenance operation uses the local row buffers in a subarray, and thus having a subarray shared by multiple lock regions will cause a conflict when accessing a row in a different lock region that maps to the subarray under maintenance. We design and evaluate SMD assuming a default lock region size of 512-row subarrays.
Modern DRAM chips typically use the density-optimized open-bitline architecture [84, 64], which places sense amplifiers on both top and bottom sides of a subarray and adjacent subarrays share sense amplifiers. With the open-bitline architecture, the MC should not access a row in a subarray adjacent to one under maintenance. To achieve this, the Lock Controller simply sends an ACT_NACK when the MC attempts to activate a row in a subarray adjacent to a locked region. Consequently, when SMD locks a region that spans 512-row subarrays, it prevents the MC from accessing subarrays in a bank with rows ( subarrays). In the folded-bitline architecture [84, 64], adjacent subarrays do not share sense amplifiers with each other. Therefore, SMD prevents access only to the subarrays in a locked region. We evaluate SMD using the density-optimized open-bitline architecture.
Ensuring Timely Maintenance Operations. A maintenance mechanism cannot lock a region with an active row. To lock the region, the maintenance mechanism waits for the MC to precharge the row. DRAM standards specify the maximum allowed time for a row to remain active as 9x tREFI [78, 83, 85]). As such, a maintenance mechanism is delayed at most by 9x tREFI.
@@ 4.3 Controlling an SMD Chip
While introducing changes in the memory controller (MC) to handle ACT_NACK, SMD also simplifies MC design and operation as the MC no longer implements control logic to periodically issue DRAM maintenance commands. For example, the MC does not 1) prepare a bank for refresh by precharging it, 2) implement timing parameters relevant to refresh, and 3) issue REF commands. The MC still maintains the bank state information (e.g., whether and which row is open in a bank) and respects the DRAM timing parameters associated with commands for performing access (e.g., ACT, PRE, RD, WR).
Handling an ACT_NACK Signal. Fig. 3 depicts a timeline that shows how the MC handles an ACT_NACK signal. Upon receiving an ACT_NACK, the MC waits for ACT Retry Interval (ARI) time to re-issue the same ACT. It keeps re-issuing the ACT command once every ARI until the DRAM chip accepts the ACT. To ensure that a DRAM chip cannot lock a region for prolonged periods, SMD enforces a minimum delay of ARI between the end of a maintenance operation and the start of the next maintenance operation targeting the same region. This allows the MC to serve a request that targets the locked region immediately after the maintenance operation in this region ends. While waiting for ARI, the MC can activate a row from a different lock region or bank, to overlap the ARI latency with a useful operation.
@@
We estimate the overall latency based on the latency of an RD in a conventional DRAM chip. An RD has a latency breakdown that resembles the latency breakdown of a . An RD command 1) propagates from the MC to the DRAM chip, 2) accesses data in a portion of the row buffer in the corresponding bank, and 3) sends the data back to the MC. In the DDR4 standard [78], the latency between issuing an RD and the first data beat appearing on the data bus is defined as tCL (typically 22 cycles for DDR4-3200). The latency components 1) and 3) of RD are similar to those of . Thus, the main difference between and tCL arises from the second component. According to our evaluation, the latency of accessing the Lock Region Table (LRT) is (section 7.4). Given the relatively low complexity of the LRT compared to the datapath that is involved during an RD, we believe the overall latency can be designed to be much smaller than tCL. We assume cycles unless stated otherwise. In our evaluations, we find that small latencies (e.g., ) have negligible effect on system performance mainly because the number of rejected ACTs constitute a small portion of all ACTs.
4.4 ACT_NACK Divergence Across DRAM Chips
SMD maintenance operations take place independently in each DRAM chip. Therefore, when a DRAM rank, which can be composed of multiple DRAM chips operating in lock step, receives an ACT command, some of the chips may send an ACT_NACK while others do not. Normally, ACT_NACK divergence would not happen for maintenance mechanisms that perform the exact same operation at the exact same time in all DRAM chips (e.g., the Fixed Rate Refresh mechanism, Section 5.1). However, a mechanism can also operate differently in each DRAM chip, e.g., a variable rate refresh mechanism [17, 15, 38]. As a result of this divergence, the row becomes partially open in chips that do not send ACT_NACK.
As a solution, the MC waits for ARI and re-issues the same ACT command to retry activating the row in the chips that previously sent an ACT_NACK signal. 555SMD retries activating the DRAM row in chips that rejected the previous activate command by sending a new activate command to all DRAM chips in the rank. DRAM chips that already have an active row simply ignore the activate command. Our analysis shows that under the worse-case ACT_NACK divergence (i.e., when each of the 16 chips in a rank perform refresh on a different region at different time), SMD-VR performs 15.4% worse than SMD-VR with no ACT_NACK divergence, which still provides 7.8% average speedup over conventional DDR4 refresh for memory-intensive four-core workloads.
4.5 Ensuring Forward Progress for Memory Requests
@@
The MC could fail to retry an ACT_NACK’d memory request if an active DRAM row (e.g., in a lock region other than the one targeted by the rejected memory request) precludes retrying the ACT_NACK’d memory request every ARI. Such retry failures could occur repeatedly and lead to temporary starvation of a lock region (i.e., no request targeting that lock region is served for a prohibitively long time). Temporary starvation of a lock region is very unlikely to happen because it requires a maintenance mechanism to repeatedly lock the same lock region for maintenance, which is not a property of a carefully designed maintenance mechanism. Instead, a carefully designed maintenance mechanism alternates between different lock regions every as it performs new maintenance operations (e.g., periodic refresh section 5.1).
@@
To prevent temporary starvation of lock regions irrespective of maintenance mechanism design, SMD issues an ACT_NACK’d memory request every ARI. No memory request is ACT_NACK’d by two separate maintenance operations because an SMD chip does not perform any maintenance operations for ARI after the end of a maintenance operation in a lock region (section 4.3). Issuing a rejected memory request strictly every ARI does not prevent the MC from serving memory requests that access other lock regions. Within one ARI () the MC can 1) activate a DRAM row other than the ACT_NACK’d DRAM row, 2) serve seven (based on the standard values of , , timing parameters) READ/WRITE requests to this newly activated row, and 3) precharge the bank such that an activate command to the previously ACT_NACK’d DRAM row can be issued every ARI.
@@
Theorem. If the MC retries every ARI then makes forward progress.
@@
Memory Request Stall Time. To limit the maximum potential memory request stall time, 1) long time taking maintenance operations can be divided into many smaller short time taking maintenance operations, 2) the maximum time a maintenance operation takes can be specified in the DRAM standard. In our design, the most time taking maintenance operation is ECC scrubbing (Section 5.3), which takes to read a DRAM row. This stall time is comparable to what a memory request in a modern DDR4-based system experiences. In such a system, a request that arrives at the memory controller’s request queues immediately after the memory controller schedules a periodic refresh operation (by issuing a command) has to wait for the refresh operation to complete (e.g., for [78]).
@@ 4.6 Impact to Request Scheduling
The key drawback of our SMD implementation is that it makes the synchronous DDRx interface less predictable. However, the practical performance overheads of this drawback are very low, as we demonstrate in Section 7.
In this paper, we comprehensively describe only one out of many different possible implementations of SMD due to page limits. This implementation prioritizes simplicity. However, the DRAM standard can more strictly specify exactly when and for how long SMD chips are allowed to perform maintenance operations. This stricter implementation would make SMD chips as predictable as DDRx chips today. We hope that future work building on SMD will comprehensively address these issues. To make it more evident that SMD can be implemented in a way that preserves the predictability of the DDRx interface from the perspective of the memory controller (i.e., in a way that keeps the DRAM interface synchronous), we describe a (more) predictable SMD implementation.
Designing a Predictable SMD Interface. A simple way of preserving the DDRx interface’s synchronicity with SMD is to allow the DRAM chip to perform a maintenance operation periodically only at well-defined time intervals. In other words, the DRAM chip can respond with ACT_NACKs only for the fraction of time during which the chip is allowed to execute maintenance operations. We call this fraction of time the maintenance operation time (MOT). Outside of MOT, the DDRx operates as defined today (i.e., synchronously and predictably). During an MOT, the memory controller is allowed to access DRAM cells at the cost of potentially getting ACT_NACK’d by the DRAM chip. The new SMD standard would specify i) how long an MOT is, ii) the period at which the interface enters an MOT, iii) how entry to an MOT is triggered (e.g., by a new MOT command), and iv) how many MOTs the memory controller can postpone.

### v7

@@ 3.1 DRAM Refresh
DDR5 introduces Same Bank Refresh (SBR), which refreshes one bank in each bank group at a time instead of refreshing all banks concurrently as in DDR4 [141, 153]. SBR improves bank availability as the MC can access the non-refreshing banks while other banks are being refreshed. DDR5 implements SBR with a new REFsb command [119]. Implementing REFsb necessitates changes in the DRAM interface.
@@ 3.2 RowHammer Protection
In DDR4, DRAM vendors implement in-DRAM RowHammer protection mechanisms by performing Target Row Refresh (TRR) [64, 66] within the time slack available when performing regular refresh. Prior works show that TRR is vulnerable to certain memory access patterns [66, 64, 138]. DDR5 specifies Refresh Management (RFM) [120, 119] that an MC implements to aid in-DRAM RowHammer protection. As part of RFM, the MC uses counters to keep track of row activations to each DRAM bank. When a counter reaches a specified threshold value, the MC issues the new RFM command to DRAM chips. A DRAM chip then performs an undocumented operation (one that may or may not refresh one or more victim rows based on an unknown victim row selection criteria) with the time allowed for RFM to mitigate RowHammer. Implementing RFM requires significant changes in the DRAM interface and the MC design.
@@ 3.3 Memory Scrubbing
DDR5 adds support for on-die ECC [10] and in-DRAM scrubbing [119]. A DDR5 chip internally performs ECC encoding and decoding when the chip is accessed. To perform DRAM scrubbing, the MC must periodically issue the new scrub command (for manual scrubbing) or all-bank refresh command (for automatic scrubbing). Similar to Same Bank Refresh and RFM, enabling in-DRAM scrubbing necessitates changes in the DRAM interface and in the MC design.
Our goal is to accelerate the process of introducing new in-DRAM maintenance operations and enabling more efficient maintenance operations.
4 Self-Managing DRAM (SMD)
We propose the Self-Managing DRAM (SMD) framework that decouples the responsibilities of the DRAM chip and the memory controller with a single change to the DRAM interface. Giving a DRAM chip the ability (i.e., the breathing room) to autonomously perform maintenance operations renders DRAM interface modifications for implementing future maintenance operations unnecessary. SMD is part of a continuum between fully master-slave (completely controlled by one side) and fully request-reply (equal partner) interfaces. Today’s commodity DRAM interface (e.g., DDRx [119, 140, 141], HBMx [143, 144]) is completely on the fully master-slave end of the continuum as the memory controller dictates everything the DRAM chip does. SMD, in contrast, tries to achieve a balance that enables faster innovation and easier adoption of new ideas by giving the DRAM chip some freedom to autonomously perform maintenance operations. We believe that doing so would ultimately more quickly enable better (more performant, efficient, robust) computing systems.
@@ 4.1 Overview of SMD
Organization. SMD preserves the general DRAM interface and uses a single uni-directional pin (which already exists in DDR4/5, see section 6) in the physical interface of a DRAM chip. This pin is used for sending the ACT_NACK signal from the DRAM chip to the MC.
Fig. 2 shows the organization of a bank in an SMD chip. SMD divides a DRAM bank into multiple lock regions of equal size. Each bank has at least one lock region. SMD stores an entry in a per-bank Lock Region Bitvector (LRB) to indicate whether or not a lock region is reserved for performing a maintenance operation.
Operation. A maintenance operation takes place as the Lock Controller sets the lock bit in the LRB entry that corresponds to the lock region on which the maintenance operation is to be performed. When the MC attempts to open a row in a locked region, the Lock Controller generates an ACT_NACK. When a maintenance operation attempts to lock a region 1) with an open row or 2) that is already locked by another maintenance operation, the Lock Controller responds with cannot lock.
A maintenance operation and an ACT can be performed concurrently in the same bank on different lock regions.555An ECC scrubbing operation (section 5.3) uses the data bus, which is a shared resource across all lock regions. Therefore, a scrubbing operation reserves all lock regions inside a bank at once to prevent any other maintenance operation or memory access from happening at the same time. To enable this, SMD implements a row address latch in order to drive two local row decoders with two independent row addresses.666One row address latch per lock region enables two or more maintenance operations to concurrently happen on different lock regions within a bank. To keep our design simple, we restrict a maintenance operation to happen in one lock region while the MC accesses another lock region. Our design can be easily extended to increase concurrency of maintenance operations or accesses across multiple lock regions, similarly to SALP [148]. When a maintenance operation and an ACT arrive at the same lock region at the same time, the bank prioritizes the maintenance operation and the SMD chip rejects the ACT.
@@ 4.2 Region Locking Mechanism
A maintenance operation can be performed on a lock region only while the region is locked. Therefore, a maintenance mechanism must lock the region that includes the rows that should undergo maintenance. A maintenance mechanism can only lock a region that is not locked, which prevents different maintenance mechanisms from interfering with each other.
Lock Region Size. A lock region consists of a fixed number of consecutively-addressed DRAM rows. To simplify the design, we set the lock region size to one or multiple subarrays. This is because a maintenance operation uses the local row buffer in a subarray. We design and evaluate SMD assuming a default lock region size of subarrays. We evaluate SMD’s sensitivity to lock region size (for a DRAM bank with 256 subarrays) and find that increasing the number of lock regions in a bank beyond 16 (i.e., having fewer than 16 subarrays in a lock region) yields diminishing returns (section 9.2).
Modern DRAM chips typically use the density-optimized open-bitline architecture [154, 155, 156, 157], which places sense amplifiers on both sides of a subarray and where adjacent subarrays share sense amplifiers. With the open-bitline architecture, the MC should not access a row in a subarray adjacent to one under maintenance. To achieve this, the Lock Controller simply sends an ACT_NACK when the MC attempts to activate a row in a subarray adjacent to a locked region. Consequently, when SMD locks a region that spans subarrays, it prevents the MC from accessing subarrays. We evaluate SMD using the density-optimized open-bitline architecture. In the folded-bitline architecture [154, 156, 155], adjacent subarrays do not share sense amplifiers with each other. Therefore, SMD prevents access only to the subarrays in a locked region.
Ensuring Timely Maintenance Operations. A maintenance mechanism cannot lock a region with an active row. To lock the region, the maintenance mechanism waits for the MC to precharge the row. DRAM standards specify the maximum allowed time for a row to remain active as 9x tREFI [141, 153, 158, 159]). As such, a maintenance mechanism is delayed at most by 9x tREFI.
@@ 4.3 Controlling an SMD Chip
While introducing changes in the memory controller (MC) to handle ACT_NACK, SMD also simplifies MC design and operation as the MC no longer implements control logic to issue DRAM maintenance commands. For example, the MC does not 1) prepare a bank for refresh by precharging it, 2) implement timing parameters relevant to refresh, and 3) issue REF commands. The MC still maintains the bank state information (e.g., which row is open in a bank) and respects the DRAM timing parameters associated with other DRAM commands (e.g., ACT).
Handling an ACT_NACK Signal. Fig. 3 depicts a timeline that shows how the MC handles an ACT_NACK signal. Upon receiving an ACT_NACK, the MC waits for ACT Retry Interval (ARI) time to re-issue the same ACT. It can keep re-issuing the ACT command once every ARI until the DRAM chip accepts the ACT. To ensure that a DRAM chip cannot lock a region for prolonged periods, SMD enforces a minimum (configurable) delay of ARI between the end of a maintenance operation and the start of the next maintenance operation targeting the same region. This allows the MC to serve a request that targets the locked region immediately after the maintenance operation in this region ends. While waiting for ARI, the MC can activate a row from a different lock region or bank, to overlap the ARI latency with a useful operation. An enforced minimum delay of ARI is sufficient to guarantee forward progress for memory requests (section 4.4). The system designer can set the enforced minimum delay to a multiple of ARI to enforce a larger time period between two maintenance operations targeting the same region.
@@
We estimate the overall latency based on the latency of RD (read) in a conventional DRAM chip. RD has a latency breakdown that resembles the latency breakdown of a . RD command 1) propagates from the MC to the DRAM chip, 2) accesses data in a portion of the row buffer in the corresponding bank, and 3) sends the data back to the MC. In the DDR4 standard [141], the latency between issuing an RD and the first data beat appearing on the data bus is defined as tCL (typically 22 cycles for DDR4-3200). The latency components 1) and 3) of RD are similar to those of . Thus, the main difference between and tCL arises from the second component. According to our evaluation, the latency of accessing the Lock Region Bitvector (LRB) is (section 6). Given the relatively low complexity of the LRB compared to the datapath that is involved during RD, the overall latency can be designed to be much smaller than tCL. We assume cycles unless stated otherwise. In our evaluations (see section 8), we find that small latencies (e.g., ) have a negligible effect on system performance mainly because the number of rejected ACTs constitutes a small portion of all ACTs.
4.4 Forward Progress for Memory Requests
@@
The MC could fail to retry an ACT_NACK’d memory request if an active DRAM row (other than the one accessed by ACT_NACK’d memory request) prevents retrying the ACT_NACK’d memory request every ARI. Such retry failures could occur repeatedly and lead to temporary starvation of a lock region (i.e., no request targeting that lock region is served for a prohibitively long time). Temporary starvation of a lock region is very unlikely to happen because it requires a maintenance mechanism to repeatedly lock the same lock region for maintenance, which is exactly what a carefully designed maintenance mechanism should avoid. Instead, a carefully designed maintenance mechanism alternates between different lock regions as it performs new maintenance operations (e.g., periodic refresh section 5.1).
@@
To prevent temporary starvation of lock regions irrespective of maintenance mechanism design, an implementation of SMD issues an ACT_NACK’d memory request every ARI. No memory request is ACT_NACK’d by two separate maintenance operations because an SMD chip does not perform any maintenance operations for at least ARI after the end of a maintenance operation in a lock region (section 4.3). Issuing a rejected memory request strictly every ARI does not prevent the MC from serving memory requests that access other lock regions.777An SMD chip may reject a memory request’s activate command. This means that the memory scheduler, based on its scheduling algorithm, has deemed that request to be the most important request to schedule. Therefore, it should be in the best interest of the scheduling algorithm to serve this memory request as soon as possible. The ARI interval allows the memory controller to schedule one or multiple other memory requests by leveraging lock-region-level parallelism. Within one ARI () the MC can 1) activate a DRAM row other than the ACT_NACK’d DRAM row, 2) serve seven (based on the DDR4 standard values of , , timing parameters) READ/WRITE requests to this newly activated row, and 3) precharge the bank such that an activate command to the previously ACT_NACK’d DRAM row can be issued every ARI. To ensure that the memory controller retries the rejected memory request in due time, we set the number of allowed column commands (RD/WR) to an open row before an older request (to another row) is serviced (i.e., the Cap parameter [160]). Based on ARI=, we use a Cap of 7 in our evaluations.
@@
Theorem. If the MC retries every ARI then makes forward progress.888 The MC is not too busy to retry the request because we modify the page policy by capping the number of column accesses in an ARI to 7.
@@
Memory Request Stall Time. To limit the maximum potential memory request stall time, 1) time-intensive maintenance operations can be divided into many small, short-duration maintenance operations, 2) the maximum time a maintenance operation takes can be specified in the DRAM standard. In our design, the longest-duration maintenance operation is ECC scrubbing (Section 5.3), which takes to read a DRAM row (assuming no corrected ECC codewords need to be written back to the DRAM array). This stall time is comparable to what a memory request in a modern DDR4-based system could experience. In such a system, a request that arrives at the memory controller’s request queues immediately after the memory controller schedules a periodic refresh operation (by issuing a command) has to wait for the refresh operation to complete (e.g., for [141]).
@@ 4.5 Impact to Request Scheduling
The key drawback of our SMD implementation is that it makes the synchronous DDRx interface less predictable (i.e., slightly asynchronous) from the perspective of the MC. However, the practically-observed performance overheads of this drawback are very low, as we demonstrate in section 8.
In this paper, we comprehensively describe only one of many different possible implementations of SMD due to page limits. This implementation prioritizes simplicity. However, the DRAM standard can more strictly specify exactly when and for how long SMD chips are allowed to perform maintenance operations. This stricter implementation would likely make SMD chips as predictable as DDRx chips today. We hope that future work building on SMD can comprehensively study such issues. To make it more evident that SMD can be implemented in a way that preserves the predictability of the DDRx interface from the perspective of the memory controller (i.e., in a way that keeps the DRAM interface synchronous), we briefly describe a (more) predictable SMD implementation.
Designing a Predictable SMD Interface. A simple way of preserving the DDRx interface’s synchronicity with SMD is to allow the DRAM chip to perform a maintenance operation periodically only at well-defined time intervals. In other words, the DRAM chip can respond with ACT_NACKs only for the fraction of time during a mode in which the chip is allowed to execute maintenance operations. We call this mode the maintenance operation time (MOT). Outside MOT, the DDRx operates as defined today (i.e., synchronously and predictably). During MOT, the memory controller is allowed to access DRAM cells at the cost of potentially getting ACT_NACK’d by the DRAM chip. The new SMD standard would specify i) how long MOT is, ii) the period at which the interface enters MOT (MOT_period), iii) how entry to MOT is triggered (e.g., by a new MOT command), and iv) how many MOTs the memory controller can postpone.